PAYMENT AUTHORIZATION FORM

I hereby authorize China International Conference Center for Science and Technology (CICCST) to debit the below mentioned credit card account with the total value of the items booked by me stated in this form.

Total Value to Be Debited: US$

For: registration fee of ICM 2002 US$ hotel cost/deposit US$ tours/others US$

The payment is for (name(s) of delegate(s))

Card Name: □VISA □Diners Club □MasterCard □American Express □JCB

Card Number:

Expiration Date (year/month/date)

Name of Card Holder:

Card Holder’s Organization/Address:

Card Holder’s Fax/Email (in case problem in payment arises):

Card Holder’s Signature: Date of Signing:

Note: Sometimes credit card companies reject our request for payment; some place a limit on the amount of payment. In both cases, we will inform you as soon as possible.

Please FAX this form to (+86-10) 6218 0142, attention: Mr. Liu Feng, CICCST, or mail toMr. Liu Feng, ICM2002, CICCST, 86 Xueyuan Nan Road, Beijing 100081, China

International Congress of Mathematicians, Beijing, China, August 20-28, 2002